# Simultaneous Inference in Multiple Matrix-Variate Graphs for High-Dimensional Neural Recordings

### v1

@@ 5.2 Experimental Data Analysis
We analyzed local field potential recordings (LFPs) from two Utah arrays implanted in the prefrontal cortex (PFC) and visual area V4 of a Macaque monkey (Fig. 3(a)). Each Utah array had $96$ electrodes, and the neuroelectrical activity in the two areas were simultaneously recorded during a memory-guided saccade task. One trial of the task consisted of the following stages (see Fig. 3(b)):
@@
    The monkey fixated at the center of the screen for $200$ ms.
@@
    The monkey had to remember the target location during a delay period of $500$ ms while maintaining fixation.
@@
    After the delay period, the fixation point turned off, and the monkey had to make a saccade to the remembered target location.
See Khanna et al., (2020) for the details of the experiment and data collection. In previous work in similar tasks, V4 has been reported to retain higher order information (e.g., color and shape) and attention to visual objects (Orban,, 2008; Fries et al.,, 2001), while prefrontal cortex (PFC) is considered to exert cognitive control in working memory (Miller and Cohen,, 2001). Despite their spatial distance and functional differentiation, these regions have been presumed to cooperate during visual memory retention (Sarnthein et al.,, 1998; Liebe et al.,, 2012).
Figure 3: (a) The positions of analyzed cortical areas, V4 and PFC in a primate brain. Area PFC has been associated with top-down control of attention while area V4 is a mid-level visual area with robust visual responses. The neuroelectrical activity in each area was recorded by a 96-electrode Utah array. (b) The timeline of one experimental trial (Khanna et al.,, 2020). We are interested in the fixation, target presentation and delay stages of $750$ ms in total. (c) LFP recordings for one experimental trial. Each x-axis indicates $96$ electrodes in each brain area, and the y-axis is time in ms. Time $t=0$ was aligned at the start of the delay period.
The analyzed dataset was collected over $m=5$ experimental sessions, which consisted of $n_{l}=2000,2995,3000,3000$ and $3000$ successful trials. For each trial, we observed matrix-variate data of temporal dimension $p=750$ ($750$ ms at sampling rate $1$ kHz) and spatial dimension $q=192$ ($2\times 96$ electrodes) in total. The data of one example trial is shown in Fig. 3(c). The color in the heatmap corresponds to the intensity of the LFP signal. The data are available in a KiltHub repository of Carnegie Mellon University, located at http://10.1184/R1/19248827.v1.
Our objective is to detect changes in the spatial correlation structure within and between the brain regions across four experimental stages: fixation ($200$ ms), target presentation ($50$ ms), early delay (the first $250$ ms during delay), and late delay (the last $250$ ms during delay). We applied the proposed method in Section 3 separately to each stage.
@@
First, we applied our method separately to each brain area. We set $\gamma_{i}$ to be the same value $\gamma$ and used the cross-validation method to determine the group Lasso tuning parameter $\gamma$. Using the spatial location of the electrodes, we inferred the relationship between neural connectivity and physical distance. We looked into the average of the spatial partial correlation estimates $\widehat{\rho}_{ij}^{(\mathcal{S})}=\frac{1}{\sum_{l=1}^{m}\sqrt{n_{l}p}}\sum_{l=1}^{m}\sqrt{n_{l}p}~{}\widehat{\rho}^{(\mathcal{S},l)}_{ij}$ on groups of spatially equidistant edges on the electrode array. Fig. 4 demonstrates the monotonic decreasing relationship of the averaged $\widehat{\rho}^{(S)}_{ij}$ with the spatial distance at the late delay stage in V4. The same trends were observed for the other experimental stages and in PFC, advocating the previous discoveries about the strong dependency of correlated neural activity on the physical distance (Goris et al.,, 2014; Vinci et al.,, 2018). This result provides a sanity check for our spatial partial correlation estimates.
Figure 4: The averaged $\widehat{\rho}^{(\mathcal{S})}_{ij}$ within a group of spatially equidistant within-area edges during the late delay stage in V4. Notice that the spatial partial correlation declined as the spatial distance increased. This phenomenon is consistently identified over all experimental stages and both brain areas.
@@ 5.2.2 Within-area Inference
Next, for each electrode $i$, we evaluated its overall connectivity within the same region by $\sum_{j}\lvert\widehat{\rho}_{ij}^{(\mathcal{S})}\rvert$. Fig. 5 shows the smoothed distributions of within-area connectivity for the four experimental stages and in the two brain areas. We first observed that V4 exhibited more within-area connectivity than PFC. In V4, the within-area connectivity was strongest during the fixation and cue stages and declined during the delay stages. On the other hand, the level of connectivity in PFC remained stable over the experimental stages.
@@ 5.2.3 Cross-area Inference
Last, we applied our method to the recordings from both areas and identified significant cross-area connectivity. To reduce computation time and collinearity, we subsampled the electrodes in each area by taking every other node along the spatial dimension, reducing the spatial dimension $q$ from $192$ to $50$ in total. This reduced the number of cross-area edges to $625$. Also, because beta oscillations are often associated with communication across the two brain areas (Klein et al.,, 2020; Miller et al.,, 2018), we first band-pass filtered the LFP recordings at the beta band $15-30$ Hz and downsampled the filtered data to $200$ Hz. Fig. 6 shows the cross-area edges, color-coded by the estimated spatial partial correlations $\widehat{\rho}^{(S)}_{ij}$. We found that the cross-area connectivity between V4 and PFC was much stronger between some electrode pairs than others, and changed substantially over time. In particular, the edges with electrode $e_{1}$ (shown red in Fig. 6) strengthened over time, whereas the other edges were strongest during the target presentation. This is broadly consistent with the finding that there are interactions between V4 and PFC during attention tasks (Squire et al.,, 2013; Snyder et al.,, 2021) and that corticocortical connections at a distance are quite sparse and highly variable in strength depending on distance and response properties (Ts’o et al.,, 1986; El-Shamayleh et al.,, 2013; Smith and Kohn,, 2008).
@@
Figure 6: Estimated cross-area connectivity between PFC and V4, shown as edges color-coded by $\widehat{\rho}_{ij}^{(\mathcal{S})}$ over the four experimental stages. The x- and y-axes are spatial coordinates of the electrodes on each array. The plot shows a spatial distinction in the evolution of the cross-area connection.
We analyzed cross-area connectivity in the two spatially-distinct edge sets using $\lVert\widehat{\rho}_{E_{1}}\rVert_{\infty}$ and $\lVert\widehat{\rho}_{E_{2}}\rVert_{\infty}$, where $E_{1}$ is the cross-area edges without electrode $e_{1}$, and $E_{2}$ consists of the other edges. Fig. 7 shows the change of cross-area connectivity in the two edge sets over the four experimental stages. In $E_{1}$, we identified that the cross-area connectivity was strongest during the target presentation. This might relate to the communication between V4 and PFC about the target stimulus. On the other hand, $E_{2}$ showed enhanced connectivity during the delay stage, when the animal had to process the visual signal and prepare the subsequent decision. The monotonicity of the cross-area connectivity in edge set $E_{2}$ was tested by the linear trend test. In the test, we linear-regressed the level of connectivity at the four stages against the respective time points on the experimental timeline. If the slope of the regression function was positive, the connectivity was increasing across the stages; if negative, the connectivity was decreasing. The p-value was estimated using $100,000$ bootstrap samples. We note that this linear trend test assumed the statistics $\lVert\widehat{\rho}_{E_{2}}\rVert_{\infty}$ were independent across the stages, which might not be true. The test showed that the connectivity levels in $E_{2}$ were significantly increasing. A similar pattern was found in the averaged partial correlation in $E_{2}$ (Fig. A.1 ). This finding makes a remarkable contrast with the suppressed within-area connectivity in V4 during the delay stages, visible in Fig. 5. When we applied the same test to the averaged partial correlations within V4, the within-area connectivity in V4 was found to be significantly decreasing (Fig. A.2 ).
@@
The cross-area connectivity trend was further evidenced by the $c$-level testing. By testing if $\lVert\rho_{E_{2}}\rVert_{\infty}$ exceeds threshold level $c$, i.e.,
@@
we compared the cross-area connectivity strengths in the four stages. Table 2 shows the test results at four levels $c=0,0.02,0.03,0.04$. At $c=0$, the cross-area connectivity in $E_{1}$ was significant with $\alpha=0.05$ at all experimental stages. As the level increased, the tests at the early stages began unrejected. At $c=0.04$, the test was significant only at the late delay stage. These results are consistent with our finding in Fig. 7 and the subsequent linear trend test.
Our inference results support the previous studies that neural variability in the spiking of neurons declines during the stimulus onset (Churchland et al.,, 2010), and visual stimuli cause a substantial decrease in the correlation of cortical neurons (Smith and Kohn,, 2008). We also discovered robust sustained within-area connectivity in PFC during the delay stage, compared to V4, which was also reported by Leavitt et al., (2017).
Table 2: The $c$-level test results for $\lVert\widehat{\rho}_{E_{2}}\rVert_{\infty}$ at $c$-levels $\{0,0.02,0.03,0.04\}$. Entries with * represent the significance at $\alpha=0.05$.

### v2

@@ 5.2 Experimental Data Analysis
We analyzed local field potential recordings (LFPs) from two Utah arrays implanted in the prefrontal cortex (PFC) and visual area V4 of an experimetnal subject (a Macaque monkey, Fig. 3(a)). Each Utah array had $96$ electrodes, and the neuroelectrical activity in the two areas were simultaneously recorded during a memory-guided saccade task. One experimental trial of the task consisted of the following stages (see Fig. 3(b)):
@@
    The subject fixated at the center of the screen for $200$ ms.
@@
    The subject had to remember the target location during a delay period of $500$ ms while maintaining fixation.
@@
    After the delay period, the fixation point turned off, and the subject had to make a saccade to the remembered target location.
See Khanna et al. (2020) for the details of the experiment and data collection. Previous studies in similar tasks have shown that area V4 encodes higher order visual features such as color and shape and is modulated by visual attention (Orban, 2008; Fries et al., 2001), while the prefrontal cortex (PFC) plays a central role in cognitive control and working memory (Miller and Cohen, 2001). Despite their anatomical separation and distinct functional roles, coordinated activity between V4 and PFC has been reported during visual memory retention (Sarnthein et al., 1998; Liebe et al., 2012). Building on these findings, our objective is to examine how the spatial correlation structure within and between these regions evolves across four experimental stages: fixation ($200$ ms), target presentation ($50$ ms), early delay (the first $250$ ms of the delay period), and late delay (the last $250$ ms of the delay period).
Figure 3: (a) The positions of the analyzed cortical areas, V4 and PFC in a primate brain. The neuroelectrical activity in each area was recorded by a 96-electrode Utah array. (b) The timeline of one experimental trial. (c) LFP recordings for one experimental trial. Each x-axis indicates $96$ electrodes in each brain area, and the y-axis is time in ms. Time $t=0$ was aligned at the start of the delay period.
The dataset was collected across $m=5$ experimental sessions indexed by $l$, containing $n_{l}=2000,2995,3000,3000,$ and $3000$ successful trials, respectively. For each trial, we observed matrix-variate data with temporal dimension $p=750$, corresponding to $750$ ms recorded at a sampling rate of $1$ kHz, and spatial dimension $q=192$, corresponding to $2\times 96$ electrodes. An example trial is shown in Fig. 3(c), where the heatmap color indicates the intensity of the local field potential signal. Although the strength of partial correlations may vary across experimental sessions due to session-specific variability, we expect the underlying sparsity pattern of the connectivity graph to be largely shared across sessions within each experimental stage. To exploit this structure, we applied the proposed method in Section 3 separately to each stage, jointly fitting multiple matrix-variate graphs across sessions by treating trials within each session as a single sample group. This joint modeling strategy allows us to borrow information across sessions, thereby improving statistical power for detecting common connectivity patterns while accommodating heterogeneity in correlation magnitudes.
@@
As an initial sanity check, we examined the relationship between neural connectivity and physical distance. Previous studies have shown that LFP signals from neighboring electrodes in monkey brains tend to exhibit strong correlations and that the correlation decays sharply as the distance between electrodes increases (Leopold and Logothetis, 2003). To quantify this effect, we applied our method to the raw LFP signals from each brain region at each experimental stage. For each electrode pair $(i,j)$, we aggregated the spatial partial correlation estimates across sessions by calculating $\widehat{\rho}_{ij}^{(\mathcal{S})}:=\frac{1}{\sum_{l=1}^{m}\sqrt{n_{l}}}\sum_{l=1}^{m}\sqrt{n_{l}}~\widehat{\rho}^{(\mathcal{S},l)}_{ij}$, where the weight $\sqrt{n_{l}}$ accounts for the varying number of trials across sessions. We then grouped electrode pairs by spatial distance and calculated the average of $\widehat{\rho}_{ij}^{(\mathcal{S})}$ within each distance group. Fig. 4 shows the average spatial partial correlations within each equidistant group in area V4 during the late delay stage, plotted against physical distance. The results demonstrate a clear monotonic decrease in average $\widehat{\rho}^{(\mathcal{S})}_{ij}$ with increasing distance, consistent with prior findings. We observed similar trends in other stages and in the PFC.
Figure 4: The averaged $\widehat{\rho}^{(\mathcal{S})}_{ij}$ within a group of spatially equidistant within-area edges during the late delay stage in V4.
@@ 5.2.2 Within-area Inference
Next, we examined how within-region connectivity changes across the experimental stages. Using the same output from our method applied to each brain region at each stage, we quantified the within-region average connectivity of each electrode $i$ by computing $\sum_{j}\lvert\widehat{\rho}_{ij}^{(\mathcal{S})}\rvert$, where $\sum_{j}$ denotes summation over all other electrodes within the same region as $i$.
Fig. 5 presents the smoothed distributions of within-area connectivity across the four experimental stages in both brain regions. We observed that V4 exhibited higher within-area connectivity than PFC. In V4, connectivity was strongest during the fixation and cue stages, then declined during the delay stages. In contrast, the level of connectivity in PFC remained relatively stable across experimental stages.
@@ 5.2.3 Cross-area Inference
We applied our method to the recordings from both areas and identified significant cross-area connectivity. To reduce computation time and collinearity, we subsampled the electrodes in each area by taking every other node along the spatial dimension, reducing the spatial dimension $q$ from $192$ to $50$ in total. This reduced the number of cross-area edges to $625$. Also, because beta oscillations are often associated with communication across the two brain areas (Klein et al., 2020; Miller et al., 2018), we first band-pass filtered the LFP recordings at the beta band 153015-30 Hz and downsampled the filtered data to $200$ Hz. Fig. 6 shows the cross-area edges, color-coded by the estimated spatial partial correlations $\widehat{\rho}^{(S)}_{ij}$ between electrode $i$ in V4 and electrode $j$ in PFC. We find that cross-area connectivity between V4 and PFC is markedly heterogeneous across electrode pairs and varies substantially across experimental stages. In particular, edges involving electrode $e_{1}$ (shown in red in Fig. 6) are consistently strong across stages and further strengthen during the late delay period, when working memory demands are highest. This pattern suggests that connectivity through electrode $e_{1}$ may reflect attentional signals directed by PFC during visual memory retention. In contrast, other cross-area edges are strongest during the target presentation stage, which may reflect transient top-down signaling from V4 to PFC associated with stimulus processing. These observations are broadly consistent with prior findings that V4–PFC interactions are engaged during attention tasks (Squire et al., 2013; Snyder et al., 2021), and that long-range corticocortical connections are sparse and exhibit substantial variability in strength depending on anatomical distance and functional properties (Ts’o et al., 1986; El-Shamayleh et al., 2013; Smith and Kohn, 2008).
@@
Figure 6: Estimated cross-area connectivity between PFC and V4, shown as edges color-coded by partial correlations $\widehat{\rho}_{ij}^{(\mathcal{S})}$ between electrode $i$ in V4 and electrode $j$ in PFC over the four experimental stages. The x- and y-axes are spatial coordinates of the electrodes on each array. The plot shows a spatial distinction in the evolution of the cross-area connection.
Based on this observation, we further analyzed the cross-area connectivity in two distinct edge sets: (i) edges incident to $e_{1}$ and (ii) edges excluding $e_{1}$. Cross-area connectivity was quantified by the maximal partial correlations within each edge set across the four experimental stages. For statistical inference, we constructed 95%-confidence intervals for the maximal partial correlation in each edge set and stage, as defined in Eq. 8, using 100,000 Gaussian bootstrap samples, i.e., $\|\widehat{T}_{E}+\widehat{Z}\|_{\infty}$ where $\widehat{Z}\sim\operatorname{N}(0,\widehat{S}_{EE})$.
Fig. 7 presents these confidence intervals, with points indicating the median values of the bootstrap samples $\lVert\widehat{T}_{E}+\widehat{Z}\rVert_{\infty}$ for the two edge sets across experimental stages. For edges excluding $e_{1}$, we reaffirm that cross-area connectivity was strongest during target presentation. In contrast, edges incident to $e_{1}$ exhibited increased connectivity during the delay stage, when the animal processed the visual signal and prepared for the subsequent decision.
We tested the monotonic trend in cross-area connectivity for edges incident to $e_{1}$ using a linear trend test. In this test, we regressed connectivity levels at the four stages against the respective time points on the experimental timeline. A positive slope indicated increasing connectivity, while a negative slope indicated decreasing connectivity. The uncertainty in the slope estimate was assessed by repeating the linear trend test over the Gaussian bootstrap samples $\|\widehat{T}_{E}+\widehat{Z}\|_{\infty}$. The $95\%$-confidence interval of the slope was $(5.03\mathrm{e-}5,8.17\mathrm{e-}5)$ $\text{ms}^{-1}$, indicating that connectivity in edges incident to $e_{1}$ increased significantly over time.
A similar trend was observed when using the average partial correlation, rather than the maximal partial correlations, over the edge set (Fig. B.1). This contrasts with the decline in within-area connectivity in V4 during the delay stages, as shown in Fig. 5. When we applied the same trend test to averaged partial correlations within V4, we found that within-area connectivity in V4 decreased significantly (Fig. B.2).
In summary, (a) there is strong within-area functional connectivity (based on partial correlations) for V4 that diminishes across the four stages, while there is much weaker within-area connectivity for PFC that remains stable across the stages (Fig. 5); (b) the cross-area connectivity is relatively sparse in V4 compared with PFC, except during target illumination; the signal from electrode $e_{1}$ which, across all four stages, has the largest partial correlation (among those in V4) with those from PFC, during the final stage (late delay) has very much stronger connectivity with PFC than the signals from other electrodes in V4; also its connectivity strengthens across the stages, while connectivity of the other electrodes with PFC stays roughly the same (Figs. 6 and 7); and (c) the signal from electrode $e_{1}$ does not have especially strong connectivity within V4.
Figure 7: Inference on the maximal partial correlations over the two edge sets: (triangle) the cross-area edges with electrode $e_{1}$, marked in Fig. 6 and (circle) the other cross-area edges. The points are the median values of the bootstrap samples obtained by Algorithm 1 with $100,000$ bootstrap samples, and the error bars are the 95 % bootstrap confidence intervals.
6 Conclusion
We proposed a method for jointly estimating multiple matrix-variate graphical models that leverages shared sparsity structures across sample groups, together with a unified framework for high-dimensional inference on graph edges. Estimation is carried out using group Lasso to exploit the shared sparsity structures, while inference is designed for high-dimensional settings via state-of-the-art central limit theorems for the supremum of weighted sums. The proposed approach accommodates settings in which the spatial and temporal dimensions, as well as the number of graphs, may all diverge and potentially exceed the available sample size.
Both our model and our assumptions are motivated by practical concerns in neural data analysis. In real data, we observed that the within-area connectivity and cross-area connectivity change as the animal progresses through different experimental stages. From a scientific perspective, the data analysis suggests the intriguing possibility that there is some degree of localization in the memory task-relevant neural signals connecting V4 with PFC. However, justification for a strong scientific conclusion would require additional experimental work. The findings are most interesting as illustrations of the way changing connectivity structure, across multiple stages of a task, can be determined from high-dimensional electrode recordings.
Appendix A Supplementary results for simulation study
Table A.1 presents the average empirical coverages under the same simulation setting as in Section 5.1.2 with the number of sessions fixed at $m=3$.
Table A.1: Average of empirical coverages and their standard deviations for $m=3,p=50,q=30$.
$n$ Quantile Random Hub Band
5 0.925 0.897(0.012) 0.898(0.009) 0.908(0.010) 0.908(0.009) 0.903(0.009) 0.907(0.010)
0.95 0.935(0.006) 0.932(0.010) 0.939(0.007) 0.939(0.007) 0.934(0.008) 0.939(0.007)
0.975 0.962(0.006) 0.969(0.002) 0.970(0.004) 0.971(0.004) 0.971(0.004) 0.971(0.003)
10 0.925 0.926(0.005) 0.924(0.006) 0.923(0.005) 0.923(0.006) 0.929(0.005) 0.927(0.006)
0.95 0.944(0.003) 0.949(0.004) 0.945(0.004) 0.945(0.005) 0.953(0.005) 0.950(0.005)
0.975 0.967(0.004) 0.970(0.002) 0.978(0.003) 0.977(0.005) 0.980(0.003) 0.977(0.003)
20 0.925 0.926(0.005) 0.921(0.004) 0.926(0.005) 0.926(0.006) 0.928(0.004) 0.925(0.005)
0.95 0.954(0.005) 0.944(0.003) 0.948(0.004) 0.948(0.004) 0.951(0.005) 0.943(0.005)
0.975 0.978(0.001) 0.975(0.004) 0.977(0.005) 0.975(0.004) 0.978(0.003) 0.975(0.003)
Appendix B Supplementary results for experimental data analysis
Figure B.1 presents the confidence intervals for the mean partial correlations over the two edge sets in the four experimental stages. The mean spatial partial correlation over the edges with electrode $e_{1}$ was shown to be significantly monotonic increasing by the linear trend test ($p<1\text{e-}6$).
Figure B.1: Inference on the mean partial correlations over the two edge sets: (red) the cross-areas edges with electrode $e_{1}$, marked in Fig. 6 and (black) the other cross-areas edges. The points are the median values of the bootstrap samples obtained by Algorithm 1 with $100,000$ bootstrap samples, and the error bars are the 95 % bootstrap confidence intervals.
Figure B.2 presents the confidence intervals for the mean partial correlations within region V4 in the four experimental stages. The mean spatial partial correlation within V4 was shown to be significantly monotonic decreasing by the linear trend test ($p<1\text{e-}6$).
Figure B.2: Inference on the mean partial correlations within V4. The points are the median values of the bootstrap samples obtained by Algorithm 1 with $100,000$ bootstrap samples, and the error bars are the 95 % bootstrap confidence intervals.
@@ Appendix B Supplementary results for experimental data analysis
The cross-area connectivity trend was further evidenced by the $c$-level testing. Our simultaneous testing approach of multiple edges can also be extended to $c$-level tests (Qiu and Zhou, 2020). By testing if $\lVert\rho_{E_{2}}\rVert_{\infty}$ exceeds threshold level $c$, i.e.,
@@
we compared the cross-area connectivity strengths in the four stages. Table B.1 shows the test results at four levels $c=0,0.02,0.03,0.04$. At $c=0$, the cross-area connectivity in $E_{1}$ was significant with $\alpha=0.05$ at all experimental stages. As the level increased, the tests at the early stages began unrejected. At $c=0.04$, the test was significant only at the late delay stage. These results are consistent with our finding in Fig. 7 and the subsequent linear trend test.
Table B.1: The $c$-level test results for $\lVert\widehat{\rho}_{E_{2}}\rVert_{\infty}$ at $c$-levels $\{0,0.02,0.03,0.04\}$. Entries with * represent the significance at $\alpha=0.05$.
@@
Appendix C Proofs
Throughout the following theoretical arguments, we use $X^{(\mathcal{S},l)}$ (or $X^{(\mathcal{T},l)}$) to notate stacked spatial (or temporal) observation over temporal (or spatial) components and samples within session $l$:
$\begin{split}X^{(\mathcal{S},l)}&=\left[X^{(1,l)\top},\dots,X^{(n_{l},l)\top}\right]^{\top}\in\mathbb{R}^{n_{l}p\times q}\\ X^{(\mathcal{T},l)}&=\left[X^{(1,l)},\dots,X^{(n_{l},l)}\right]^{\top}\in\mathbb{R}^{n_{l}q\times p}.\end{split}$
These notations help us with connecting matrix-variate Gaussian graphical models to node-wise regression models:
$X^{(\mathcal{S},l)}_{ti}=X^{(\mathcal{S},l)}_{t,\cdot}\beta^{(\mathcal{S},l)}_{\cdot,i}+\epsilon^{(\mathcal{S},l)}_{ti}$
for $t=1,\dots,n_{l}p$ where $X^{(\mathcal{S},l)}_{t}$ are now dependent and not identically distributed due to the temporal association, unlike the usual regression regime. Also, for any vector $b^{(\cdot)}:=\{b^{(l)}\in\mathbb{R}^{p}\}_{l=1,\dots,m}$, we use $\underline{b}^{(\cdot)}$ to denote the collection of standardized elements, i.e.,
$\underline{b}^{(\cdot)}_{j}:=\left\{\frac{\lVert X^{(\mathcal{S},l)}_{\cdot,j}\rVert_{2}}{\sqrt{n_{l}p}}b^{(l)}_{j}\right\}_{l=1,\dots,m}$
so that Eq. 2 is expressed in a more canonical form: